OpenStudy (kj4uts):

For f(x)=4x+1 and g(x)=x^2-5, find (f-g)(x) a. -x^2 + 4x - 4 b. 4x^2 - 19 c. 6 + 4x - x^2 d. x^2 - 4x - 6 I think it's c. please explain!

OpenStudy (anonymous):

oh, you are right, hehehe.. me wrong. It's C

OpenStudy (kj4uts):

I was originally getting -x^2 + 4x + 6 when I combined like terms but if you look at c. it is just that but backwards!

OpenStudy (anonymous):

So that is why I said you are correct

OpenStudy (zpupster):

(f-g) (x) = f(x) - g(x) = 4x+1 - x^2 + 5 = -x^2 + 4x +6

OpenStudy (kj4uts):

Thanks for your help both of you just wanted to confirm!